# Lipschitz-Based Robustness Certification Under Floating-Point Execution

### 17 Mar 2026 v2

@@ Abstract.
As motivating evidence, we exhibit concrete counterexamples showing that real arithmetic robustness guarantees can fail under floating-point execution, even for previously verified certifiers, with discrepancies becoming pronounced at lower-precision formats such as float16. We then develop a formal, compositional theory relating real arithmetic Lipschitz-based sensitivity bounds to the sensitivity of floating-point execution under standard rounding-error models, specialised to feed-forward neural networks with ReLU activations. We derive sound conditions for robustness under floating-point execution, including bounds on certificate degradation and sufficient conditions for the absence of overflow. We formalize the theory and its main soundness results, and implement an executable certifier based on these principles, which we empirically evaluate to demonstrate its practicality.
@@ 2.3. Verified Certified Robustness Counterexamples
We stress that this search procedure exists merely to find instances that demonstrate how, due to floating-point rounding, sensitivity-based robustness certifiers can give misleading answers—even when formally verified. It is not designed to try to produce instances in which the certified robust point $x_{0}$ is necessarily a natural input for the neural network classifier in question, nor for which the found $\varepsilon$ is necessarily semantically meaningful to the classification problem for which the neural network has been trained.
@@
Figure 1. Instances for which Tobler et al.’s verified certifier produces misleading robustness certifications. $x_{\mathit{nat}}$ is the original test point; $x_{0}$ is certified robust at $\varepsilon$, yet there exists $x_{1}$ within $\varepsilon$ of $x_{0}$ that is classified differently (due to floating-point rounding).
@@
We conclude that certifier conservatism alone does not eliminate the semantic gap between real arithmetic certification and floating-point execution. This motivates the need for a formal account of this gap and how we can address it.
@@ 6. Robustness Certification under Floating-Point Execution
The certification condition necessarily involves comparing two network executions: at the center $x$, where we observe the floating-point margin $\hat{m}_{j,i^{*}}(x)$ and check if it is sufficiently large; and at arbitrary $x^{\prime}\in B(x,\varepsilon)$, where we must guarantee $\hat{m}_{j,i^{*}}(x^{\prime})>0$. Each execution incurs floating-point rounding error. We therefore introduce two error terms, $E_{\text{ctr}}$ and $E_{\text{ball}}$, to bound the floating-point deviation in the margin at each point $x$ and $x^{\prime}$ respectively. The deviation bound at a point depends on the maximum radius at which the network operates. To distinguish the two evaluations, we make the center $x$ and perturbation radius $\varepsilon$ explicit in the notation for $r_{\ell}$ and $D_{\ell}$ (which were left implicit in Sections 4 and 5), writing e.g., $r_{\ell}(x,\varepsilon)$ and $D_{\ell}(x,\varepsilon)$:
@@
    $E_{\text{ctr}}\;:=\;\alpha_{L}^{(j,i^{*})}\cdot D_{\text{ctr}}+\beta_{L}^{(j,i^{*})}(r_{\text{ctr}}).$
@@
    $E_{\text{ball}}\;:=\;\alpha_{L}^{(j,i^{*})}\cdot D_{\text{ball}}+\beta_{L}^{(j,i^{*})}(r_{\text{ball}}).$
@@
    The certification condition holds for each $j\neq i^{*}$, where $E_{\text{ctr}}$ and $E_{\text{ball}}$ are the error bounds defined above for the pair $(i^{*},j)$:
    $\hat{m}_{j,i^{*}}(x)-L_{j,i^{*}}\cdot\varepsilon-(E_{\text{ctr}}+E_{\text{ball}})\;>\;0.$
@@
The classical real arithmetic margin certification approach (Section 2.1) requires $m_{j,i^{*}}(x)>L_{j,i^{*}}\cdot\varepsilon$ for each competing class $j$. Thus the additional term $(E_{\text{ctr}}+E_{\text{ball}})$ quantifies the margin degradation due to floating-point errors. This degradation is instance-dependent, varying with both the input point and perturbation radius, as well as the floating-point format and network complexity. Generally speaking, the margin degradation is much smaller for higher-precision formats, and worse for lower-precision ones; Section 8 quantifies the margin degradation for practical models.
@@ 7. Pre-Deployment Hybrid Certification
Moreover, this means we can instantiate Lemma 6.1 for the single input point $x$ and $\varepsilon=0$ with $D_{L-1}^{\mathrm{hybrid}}(x)$ in place of $D_{L-1}$. As a result, we can use $D_{L-1}^{\mathrm{hybrid}}(x)$ in place of $D_{\mathrm{ctr}}$ when computing $E_{\mathrm{ctr}}$ when performing the certification check of Theorem 6.2, i.e., $E_{\mathrm{ctr}}$ can be soundly replaced by:
$E_{\text{ctr}}^{\mathrm{hybrid}}\;:=\;\alpha_{L}^{(j,i^{*})}\cdot D_{L-1}^{\mathrm{hybrid}}(x)+\beta_{L}^{(j,i^{*})}(r_{\text{ctr}}).$
@@
$\hat{m}_{j,i^{*}}(x)-L_{j,i^{*}}\cdot\varepsilon-(E_{\text{ctr}}^{\mathrm{hybrid}}+E_{\text{ball}})\;>\;0.$
@@
In practice, this replacement significantly reduces the size of $E_{\mathrm{ctr}}$ and leads to more precise robustness certification, at the expense of having to run model forward passes at high precision.
Note: the same trick cannot be used for $E_{\text{ball}}$ since it captures the error for the arbitrary $x^{\prime}\in B(x,\varepsilon)$, rather than for the centre input point $x$ for which we can directly bound $x$’s deviation using the high-precision execution. For the same reason, overflow checking (Section 4) must still use the conservative quantities.
@@ 8.2.1. RQ1: Does floating-point-aware certification avoid false certificates?
In all cases, real arithmetic certifiers (including Tobler et al. (Tobler et al., 2025b)) would certify these instances as robust, despite their actual vulnerability under floating-point execution. Our floating-point-aware certifier correctly rejected all of them across all eight model-format combinations (0% certified), and even under the tighter pre-deployment hybrid certification check (Section 7), confirming that the theory successfully closes the semantic gap. By accounting for floating-point rounding throughout the network, we avoid producing false robustness guarantees for instances specifically crafted to expose the discrepancy between real and floating-point arithmetic.
@@ 8.3.1. RQ2: Is certification practical on standard benchmarks?
Table 1. Float32 certification results on standard test sets (10,000 points per model). Perturbation radii ($\varepsilon$) follow Tobler et al. (Tobler et al., 2025b). Time measurements use pre-computed weight norms. “Verified Robustness (Real)” shows percentage certified under real arithmetic assumptions (and are identical to Tobler et al. (Tobler et al., 2025b)). “Verified Robustness (FP)” shows percentage certified under floating-point execution. $\Delta$ Robustness” is the difference in percentage points between the two. “Margin Increase” shows the average relative increase in certification margin required to account for floating-point errors.
@@
Certification is practical and the theory is non-vacuous. Certification is fast (0.78–1.38ms per instance) and maintains substantial coverage (Table 1). Even accounting for floating-point effects, we certify 95.5% of MNIST instances, 82.2% of Fashion MNIST instances, and 18.9% of CIFAR-10 instances as robust at their respective perturbation radii. The gap between real and floating-point certification rates is small for MNIST (0.24 pp) and Fashion MNIST (1.49 pp) but substantial for CIFAR-10 (27.22 pp), where the larger input dimension drives greater error accumulation—reflected in the 109% average margin increase required for that model.
@@ 9. Related Work
We motivated our investigation by constructing adversarial inputs that are misleadingly certified robust (often at very small $\varepsilon$) by Lipschitz-based certifiers, despite the existence of counterexample points within $\varepsilon$ that the model classifies differently due to floating-point rounding. Jia and Rinard (Jia and Rinard, 2021) first demonstrated a similar phenomenon for complete robustness verifiers, at much larger, semantically meaningful values of $\varepsilon$. Subsequent work also showed how to construct backdoored networks together with specific trigger inputs for which a complete verifier claims robustness under its arithmetic model, yet counterexamples exist under floating-point execution (Zombori et al., 2021). More recently, Szász et al. (Szász et al., 2025) go further and show how low-level details like the order of operations, mixed-precision implementations, and other specifics of the execution runtime, can lead to robustness verifiers giving misleading results for specially crafted networks. Our theory does not explicitly handle such details, if they depart from standard floating-point semantics. We return to this point in Section 10.
@@ 10. Conclusion
Our results also shed light on the practical significance of the semantic gap between real arithmetic certification and floating-point execution. Although Lipschitz-based certification is conservative—using input-independent bounds on sensitivity that over-estimate the model’s local behaviour—our counterexamples (Section 2.3) show that this conservatism alone does not mask floating-point discrepancies—even if the observed discrepancies are small at standard precisions (e.g., float32 and above). Moreover, our floating-point-aware analysis currently has limited applicability in the lowest-precision regimes (e.g., float16) where floating-point discrepancies become largest (Section 8). These observations clarify when floating-point effects are most likely to matter for Lipschitz-based robustness certification, while our theory and implementation provide the first method for soundly accounting for these effects in Lipschitz-based robustness certification.

### 24 Mar 2026 v3

@@ Abstract.
As motivating evidence, we exhibit concrete counterexamples showing that real arithmetic robustness guarantees can fail under floating-point execution, even for previously verified certifiers. Discrepancies become pronounced at lower-precision formats such as float16, and under adversarially constructed models reach semantically meaningful perturbation radii at float32. We then develop a formal, compositional theory relating real arithmetic Lipschitz-based sensitivity bounds to the sensitivity of floating-point execution under standard rounding-error models, specialised to feed-forward neural networks with ReLU activations. We derive sound conditions for robustness under floating-point execution, including bounds on certificate degradation and sufficient conditions for the absence of overflow. We formalize the theory and its main soundness results, and implement an executable certifier based on these principles, which we empirically evaluate to demonstrate its practicality.
@@ 2.3. Verified Certified Robustness Counterexamples
We stress that, for naturally trained models, this search procedure exists merely to find instances that demonstrate how, due to floating-point rounding, sensitivity-based robustness certifiers can give misleading answers—even when formally verified. For such models it is not designed to try to produce instances in which the certified robust point $x_{0}$ is necessarily a natural input for the neural network classifier in question, nor for which the found $\varepsilon$ is necessarily semantically meaningful to the classification problem for which the neural network has been trained. As we show below, however, an adversary who controls the model weights can use the same procedure to produce counterexamples at semantically meaningful $\varepsilon$.
@@
Figure 1. Instances for which Tobler et al.’s verified certifier produces misleading robustness certifications. $x_{\mathit{nat}}$ is the original test point; $x_{0}$ is certified robust at $\varepsilon$, yet there exists $x_{1}$ within $\varepsilon$ of $x_{0}$ that is classified differently (due to floating-point rounding). Models were executed at float32.
@@
Figure 3. MNIST instances for which Tobler et al.’s verified certifier produces misleading robustness certifications, in which the float32 MNIST model has been adversarially modified to produce very large rounding errors. $x_{\mathit{nat}}$ is the original test point; $x_{0}$ is certified robust at $\varepsilon$, yet there exists $x_{1}$ within $\varepsilon$ of $x_{0}$ that is classified differently. The $\varepsilon$ values are comparable to those used for MNIST robustness evaluation throughout this paper (e.g., $\varepsilon=0.3$).
The above instances were found against models trained naturally, without any adversarial manipulation of their weights. We also investigated a stronger threat model in which the adversary controls the model weights. Starting from the MNIST model, we constructed an adversarial variant by injecting large compensating biases: a flat bias $b_{L-1}=B$ (with $B=10^{6}$) at the second-to-last hidden layer and a compensating bias $b_{L}=-(W_{L}\cdot b_{L-1})$ at the output layer, with zero biases elsewhere. In real arithmetic, the compensating bias $b_{L}$ cancels the linear contribution of $b_{L-1}$ through $W_{L}$; the cancellation is not exact because the intervening ReLU activation changes which neurons are active, but the effect on model behaviour is small. Because biases do not affect the Lipschitz constant, the real arithmetic certification is entirely unchanged. However, under floating-point arithmetic, the large bias $b_{L-1}$ inflates the intermediate values entering layer $L$, causing the matrix-vector product $W_{L}\cdot z_{L-1}$ to accumulate rounding errors proportional to these inflated values. When $b_{L}$ subsequently subtracts the large bias contribution, the small true signal is recovered but the accumulated rounding errors remain—now enormous relative to the signal. This greatly amplifies the deviation between the floating-point and real arithmetic network outputs. Crucially, the adversarial model retains 98.23% test accuracy (compared to 98.40% for the original). We applied our search procedure to this model to generate 39 counterexample instances at float32, with $\varepsilon$ values reaching up to 0.47—35 of the 39 exceeding 0.1—well within the range at which MNIST robustness is typically evaluated (e.g., $\varepsilon=0.3$ is used by Tobler et al. (Tobler et al., 2025b) and throughout this paper). Fig. 3 depicts some illustrative counterexamples. We observed the same trend for analogous adversarial models constructed for Fashion MNIST and CIFAR-10.
We conclude that certifier conservatism alone does not eliminate the semantic gap between real arithmetic certification and floating-point execution, even for naturally trained models, and that this gap can be greatly amplified by an adversary who controls the model weights—to the point of producing counterexamples at semantically meaningful perturbation radii. This motivates the need for a formal account of this gap and how we can address it.
@@ 6. Robustness Certification under Floating-Point Execution
The certification condition necessarily involves comparing two network executions: at the center $x$, where we observe the floating-point margin $\hat{m}_{j,i^{*}}(x)$ and check if it is sufficiently large; and at arbitrary $x^{\prime}\in B(x,\varepsilon)$, where we must guarantee $\hat{m}_{j,i^{*}}(x^{\prime})>0$. Each execution incurs floating-point rounding error. We therefore introduce two error terms, $E_{\text{ctr}}^{(j,i^{*})}$ and $E_{\text{ball}}^{(j,i^{*})}$, to bound the floating-point deviation in the margin at each point $x$ and $x^{\prime}$ respectively. The deviation bound at a point depends on the maximum radius at which the network operates. To distinguish the two evaluations, we make the center $x$ and perturbation radius $\varepsilon$ explicit in the notation for $r_{\ell}$ and $D_{\ell}$ (which were left implicit in Sections 4 and 5), writing e.g., $r_{\ell}(x,\varepsilon)$ and $D_{\ell}(x,\varepsilon)$:
@@
    $E_{\text{ctr}}^{(j,i^{*})}\;:=\;\alpha_{L}^{(j,i^{*})}\cdot D_{\text{ctr}}+\beta_{L}^{(j,i^{*})}(r_{\text{ctr}}).$
@@
    $E_{\text{ball}}^{(j,i^{*})}\;:=\;\alpha_{L}^{(j,i^{*})}\cdot D_{\text{ball}}+\beta_{L}^{(j,i^{*})}(r_{\text{ball}}).$
@@
    The certification condition holds for each $j\neq i^{*}$, where $E_{\text{ctr}}^{(j,i^{*})}$ and $E_{\text{ball}}^{(j,i^{*})}$ are the error bounds defined above:
    $\hat{m}_{j,i^{*}}(x)-L_{j,i^{*}}\cdot\varepsilon-(E_{\text{ctr}}^{(j,i^{*})}+E_{\text{ball}}^{(j,i^{*})})\;>\;0.$
@@
The classical real arithmetic margin certification approach (Section 2.1) requires $m_{j,i^{*}}(x)>L_{j,i^{*}}\cdot\varepsilon$ for each competing class $j$. Thus the additional term $(E_{\text{ctr}}^{(j,i^{*})}+E_{\text{ball}}^{(j,i^{*})})$ quantifies the margin degradation due to floating-point errors. This degradation is instance-dependent, varying with both the input point and perturbation radius, as well as the floating-point format and network complexity. Generally speaking, the margin degradation is much smaller for higher-precision formats, and worse for lower-precision ones; Section 8 quantifies the margin degradation for practical models.
@@ 7. Pre-Deployment Hybrid Certification
Moreover, this means we can instantiate Lemma 6.1 for the single input point $x$ and $\varepsilon=0$ with $D_{L-1}^{\mathrm{hybrid}}(x)$ in place of $D_{L-1}$. As a result, we can use $D_{L-1}^{\mathrm{hybrid}}(x)$ in place of $D_{\mathrm{ctr}}$ when computing $E_{\mathrm{ctr}}^{(j,i^{*})}$ when performing the certification check of Theorem 6.2, i.e., $E_{\mathrm{ctr}}^{(j,i^{*})}$ can be soundly replaced by:
$E_{\text{ctr}}^{\mathrm{hybrid},(j,i^{*})}\;:=\;\alpha_{L}^{(j,i^{*})}\cdot D_{L-1}^{\mathrm{hybrid}}(x)+\beta_{L}^{(j,i^{*})}(r_{\text{ctr}}).$
@@
$\hat{m}_{j,i^{*}}(x)-L_{j,i^{*}}\cdot\varepsilon-(E_{\text{ctr}}^{\mathrm{hybrid},(j,i^{*})}+E_{\text{ball}}^{(j,i^{*})})\;>\;0.$
@@
In practice, this replacement significantly reduces the size of $E_{\mathrm{ctr}}^{(j,i^{*})}$ and leads to more precise robustness certification, at the expense of having to run model forward passes at high precision.
Note: the same trick cannot be used for $E_{\text{ball}}^{(j,i^{*})}$ since it captures the error for the arbitrary $x^{\prime}\in B(x,\varepsilon)$, rather than for the centre input point $x$ for which we can directly bound $x$’s deviation using the high-precision execution. For the same reason, overflow checking (Section 4) must still use the conservative quantities.
@@ 8.2.1. RQ1: Does floating-point-aware certification avoid false certificates?
In all cases, real arithmetic certifiers (including Tobler et al. (Tobler et al., 2025b)) would certify these instances as robust, despite their actual vulnerability under floating-point execution. Our floating-point-aware certifier correctly rejected all of them across all eight model-format combinations (0% certified), and even under the tighter pre-deployment hybrid certification check (Section 7), confirming that the theory successfully closes the semantic gap. We additionally evaluated against the 39 counterexample instances generated for the adversarial biased MNIST model described in Section 2.3. At float32, our certifier correctly rejected all 39 instances under both standard and pre-deployment hybrid certification, while the real arithmetic certifier would have certified all of them. By accounting for floating-point rounding throughout the network, we avoid producing false robustness guarantees for instances specifically crafted to expose the discrepancy between real and floating-point arithmetic.
@@ 8.3.1. RQ2: Is certification practical on standard benchmarks?
Table 1. Float32 certification results on standard test sets (10,000 points per model). Perturbation radii ($\varepsilon$) follow Tobler et al. (Tobler et al., 2025b). Time measurements use pre-computed weight norms. “Verified Robustness (Real)” shows percentage certified under real arithmetic assumptions (and are identical to Tobler et al. (Tobler et al., 2025b)). “Verified Robustness (FP)” shows percentage certified under floating-point execution. $\Delta$ Robustness” is the difference in percentage points between the two. “Margin Increase” is ${\mathbb{E}_{x,j}[E_{\text{ctr}}^{(j,i^{*})}(x)+E_{\text{ball}}^{(j,i^{*})}(x)]}\big/{\mathbb{E}_{x,j}[\varepsilon\cdot L_{j,i^{*}}]}$, the average margin degradation (Section 6) relative to the average real arithmetic certification margin, where expectations are over all 10,000 test points $x$ and competing classes $j\neq i^{*}$.
@@
Certification is practical and the theory is non-vacuous. Certification is fast (0.78–1.38ms per instance) and maintains substantial coverage (Table 1). Even accounting for floating-point effects, we certify 95.5% of MNIST instances, 82.2% of Fashion MNIST instances, and 18.9% of CIFAR-10 instances as robust at their respective perturbation radii. The gap between real and floating-point certification rates is small for MNIST (0.24 pp) and Fashion MNIST (1.49 pp) but substantial for CIFAR-10 (27.22 pp), where the larger input dimension drives greater error accumulation—reflected in the 109% average margin increase required for that model. For the adversarial biased MNIST model (Section 2.3), floating-point-aware certification yields 0% verified robustness over the full 10,000-point test set under both standard and hybrid certification (compared to 95.74% under real arithmetic), confirming that the bias-induced deviation renders certification entirely vacuous at float32. The same holds for analogous adversarially biased Fashion MNIST and CIFAR-10 models.
@@ 8.4. Discussion
A similar phenomenon arises in the adversarial model setting. For the biased MNIST model (Section 2.3), certification is equally vacuous at float32: 0% of the 10,000 standard test instances are certified. However, the underlying cause is fundamentally different. For float16, the vacuousness is an inherent limitation of the lower precision format—the unit roundoff is simply too large for meaningful certification of these models. For the adversarial model, the vacuousness is a deliberate consequence of the adversary’s construction: the injected biases amplify floating-point deviation without affecting the Lipschitz constant. A certifier that ignores floating-point effects is inherently vulnerable against such a model.
@@ 9. Related Work
We motivated our investigation by constructing adversarial inputs that are misleadingly certified robust by Lipschitz-based certifiers, despite the existence of counterexample points within $\varepsilon$ that the model classifies differently due to floating-point rounding. For naturally trained models at float32 the resulting $\varepsilon$ values are very small; however, under adversarial model construction they reach semantically meaningful magnitudes (Section 2.3). Jia and Rinard (Jia and Rinard, 2021) first demonstrated a similar phenomenon for complete robustness verifiers. Subsequent work also showed how to construct backdoored networks together with specific trigger inputs for which a complete verifier claims robustness under its arithmetic model, yet counterexamples exist under floating-point execution (Zombori et al., 2021). More recently, Szász et al. (Szász et al., 2025) go further and show how low-level details like the order of operations, mixed-precision implementations, and other specifics of the execution runtime, can lead to robustness verifiers giving misleading results for specially crafted networks. Our theory does not explicitly handle such details, if they depart from standard floating-point semantics. We return to this point in Section 10.
All of the above work, like ours, exploits floating-point rounding in network execution. Jin et al. (Jin et al., 2024) target a complementary source of unsoundness: floating-point rounding in the certification computation itself. They show that for exact, conservative, and approximate certification mechanisms, rounding can cause the computed certified radius to overestimate its real-arithmetic counterpart, and develop a search method that finds adversarial examples within the inflated radius, proposing rounded interval arithmetic as a mitigation. Their attack targets the same kind of verifier-internal rounding that tools like PyRAT (Lemesle et al., 2024) address (see below). In contrast, our certifier eliminates certification-side rounding by using arbitrary-precision rational arithmetic (Section 8).
@@ 10. Conclusion
Our results also shed light on the practical significance of the semantic gap between real arithmetic certification and floating-point execution. Although Lipschitz-based certification is conservative—using input-independent bounds on sensitivity that over-estimate the model’s local behaviour—our counterexamples (Section 2.3) show that this conservatism alone does not mask floating-point discrepancies—even if the observed discrepancies are small for natural models at standard precisions (e.g., float32 and above). Furthermore, we showed that an adversary who controls the model weights can deliberately amplify these discrepancies even at float32, constructing models that retain high test accuracy yet for which a real-arithmetic certifier mistakenly certifies concrete counterexamples at semantically meaningful perturbation radii. Our floating-point-aware certifier safely rejects all such instances, though certification becomes vacuous for these adversarial models. A similar vacuousness arises in the lowest-precision regimes (e.g., float16), where floating-point discrepancies are larger and our analysis has limited applicability (Section 8). In both cases—whether caused by deliberate adversarial construction or by the lower precision format—our approach remains sound. These observations clarify when floating-point effects are most likely to matter for Lipschitz-based robustness certification, while our theory and implementation provide the first method for soundly accounting for these effects in Lipschitz-based robustness certification.